1) A square has side lengths 2 cm. What happens to its area when you triple the side lengths?
2) A cube has lengths 2 cm. What happens to its volume when you triple the side lengths?

Answer by checkley77(12844)

A=S^2
2*2=4 FOR THE AREA.
(3*2)(3*2)=6*6=36 IN THE INCREASED AREA.
--------------------------------
V=S^3
V=2^3=8 CM^3.
V=(3*2)^3=6^3=216 CM^3.